Calculate pH of ammoniac buffer consisting of 400 ml of 0.2M NH4Cl and 20 ml of 4M NH4OH. pK=4.75


1
Expert's answer
2021-03-16T08:11:16-0400

pOH = pKb + log [Salt]/[Base]

[NH4Cl] = (0.4 L * 0.2 mol/L)/(0.4 L + 0.02 L) = 0.19 mol/L

[NH4OH] = (0.02 L * 4.0 mol/L)/(0.4 L + 0.02 L) = 0.19 mol/L

pOH = 4.75 + log(0.19/0.19) = 4.75

pH = 14 - pOH = 14 - 4.75 = 9.25